OpenStudy (kva1992):

How do I get the solutions of cosx - 2sinxcosx = 0 i broght it down to cosx - sin2a = 0 but im almost positive i cant seth those equal to zero any help would be appreciated.

OpenStudy (ikram002p):

cosx - 2sinxcosx = 0 take common farctor cos x (1- 2 sin x ) = 0 right ?

OpenStudy (ikram002p):

so you would have two cases cos x =0 or 1-2 sin x =0

OpenStudy (ikram002p):

when cos x =0 then what x values (angels )that make cos x =0 ?

OpenStudy (ikram002p):

the same with 1-2 sin x =0 sin x =0.5

OpenStudy (kva1992):

wooow sometimes i just over think problems i feel soo disappointed :( but thanks for the help

OpenStudy (ikram002p):

haha np :)